OpenStudy (nuccioreggie):

Parker drove a total distance of 4 and 1 over 2 miles during the months of June and July. If Parker only drove 1 over 8 of a mile every day, which expression shows the number of days he went driving? 4 and 1 over 2 ⋅ 1 over 8 4 and 1 over 2 + 1 over 8 4 and 1 over 2 ÷ 1 over 8 4 and 1 over 2 – 1 over 8

OpenStudy (nuccioreggie):

@Michele_Laino

OpenStudy (nuccioreggie):

@tkhunny

OpenStudy (mrnood):

suppose the question was he drove 10 miles in jun and july he drove 2 miles a day how many days did he drive You can probably answer that - so you know HOW to do it just substitute the more elaborate numbers from the question into the SAME method.....

OpenStudy (tkhunny):

The idea is for you to show YOUR work. Don't post questions with no work shown.

OpenStudy (nuccioreggie):

I think its c because he dorve

OpenStudy (tkhunny):

Notational Conventions: 4 and 1 over 2 ==> 4 1/2 = 9/2 4 and 1 over 2 divided by 1 over 8 ==> (4 1/2)/(1/8) = (9/2)/(1/8) Make sure what you are writing is clear.

OpenStudy (nuccioreggie):

So wich is the correct answe

OpenStudy (nuccioreggie):

So its C correct ?? @tkhunny

OpenStudy (mrnood):

you should not ask or expect answers here - you have the method explained - try to do the work yourself...

OpenStudy (tkhunny):

You must figure out how many (1/8) there are in (9/2). That sounds like a division problem.

OpenStudy (nuccioreggie):

Yea i did i said the same thing thats why i think its C @tkhunny

OpenStudy (tkhunny):

Perfect. Move on to the next problem with a little more confidence.

OpenStudy (nuccioreggie):

Ok but do you also think that it is C as well?

OpenStudy (tkhunny):

It doesn't matter what I think. It matters what it IS. Move on to the next one.